OpenStudy (kimes82):

find the integral y√(2+10y−25y^2) dy

OpenStudy (kimes82):

wouldnt (y^2-2/5y+1/25) be to the 3/2?

zepdrix (zepdrix):

Oh there's a y in front of the integral? Did you edit the question? Or did I completely miss that..? Oh boy..

OpenStudy (kimes82):

I went the same route but i got this answer C-(1/25)(-25y^2+10y+2)^(3/2)+(1/50)(5y-1)sqrt(-25y^2+10y+2)(3/50)sin^-1((5y-1)/sqrt3))

OpenStudy (kimes82):

and haha I didnt edit it, the y was originally there

OpenStudy (kimes82):

oh wow im so confused now

Nnesha (nnesha):

WHAT is the question?!?

zepdrix (zepdrix):

This is the question:\[\large\rm \int\limits y\sqrt{2+10y-y^2}~dy\]

OpenStudy (kimes82):

yes

Nnesha (nnesha):

is that a simple question or ami sleepy?

Nnesha (nnesha):

just jooooking that's not an easy Q "D"

Nnesha (nnesha):

hmm it's -25y^2 -.p

OpenStudy (raffle_snaffle):

Can't apply U sub so integration by parts I think.

imqwerty (imqwerty):

It is of the format - \[\int(Px+Q) \sqrt{ax^2+bx+c}\]

OpenStudy (raffle_snaffle):

Let u = y and dv = sqr(stuff)

imqwerty (imqwerty):

In such cases we just put Px+Q=l(derivative of stuff under sqrt) +m where l and m are any two numbers which u can find by comparing the RHS and LHS next you write the integral like this- \(\int l(derivative ~of ~stuff~under~sqrt)+m\sqrt{ax^2+bx+c}\) And then you open the parentheses and split the integral

imqwerty (imqwerty):

Aww i forgot to put l(derivative of stuff) +m in parentheses in that integral

OpenStudy (robtobey2):

An integration using the Mathematica program is attached.